मराठी

If the letters of the word ASSASSINATION are arranged at random. Find the probability that all A’s are not coming together

Advertisements
Advertisements

प्रश्न

If the letters of the word ASSASSINATION are arranged at random. Find the probability that all A’s are not coming together

बेरीज
Advertisements

उत्तर

Total number of word is ASSASSINATION are 13.

Where, we have 3A’s, 4S’, 2I’s, 2N’s, 1T’s and 1O’s.

If all A’s are coming together, then three are 11 alphabets

Number of words when all A’s come together

= `(11!)/(4!2!2!)`

∴ Probability when all A’s come together

= `((11!)/(4!2!2!))/((13!)/(4!3!2!2!))`

= `(11!)/(4!2!2!) xx (4!3!2!2!)/(13!)`

= `(11! xx 3!)/(13!)`

= `6/(13 xx 12)`

= `1/26`

∴ Required probability when all A’s do not come together

= `1 - 1/26 = 25/26`

shaalaa.com
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 16: Probability - Exercise [पृष्ठ २९८]

APPEARS IN

एनसीईआरटी एक्झांप्लर Mathematics Exemplar [English] Class 11
पाठ 16 Probability
Exercise | Q 14.(c) | पृष्ठ २९८

व्हिडिओ ट्यूटोरियलVIEW ALL [1]

संबंधित प्रश्‍न

A die is thrown, find the probability of following events:

  1. A prime number will appear,
  2. A number greater than or equal to 3 will appear,
  3. A number less than or equal to one will appear,
  4. A number more than 6 will appear,
  5. A number less than 6 will appear.

A fair coin is tossed four times, and a person win Re 1 for each head and lose Rs 1.50 for each tail that turns up.

From the sample space calculate how many different amounts of money you can have after four tosses and the probability of having each of these amounts.


In a lottery, person chooses six different natural numbers at random from 1 to 20, and if these six numbers match with the six numbers already fixed by the lottery committee, he wins the prize. What is the probability of winning the prize in the game? [Hint: order of the numbers is not important.]


Fill in the blank in following table:

P(A) P(B) P(A ∩ B) P(A ∪ B)
0.35 ... 0.25 0.6

Fill in the blank in following table:

P(A) P(B) P(A ∩ B) P(A ∪ B)
0.5 0.35 .... 0.7

From the employees of a company, 5 persons are selected to represent them in the managing committee of the company. Particulars of five persons are as follows:

S. No. Name Sex Age in years
1. Harish M 30
2. Rohan M 33
3. Sheetal F 46
4. Alis F 28
5. Salim M 41

A person is selected at random from this group to act as a spokesperson. What is the probability that the spokesperson will be either male or over 35 years?


The number lock of a suitcase has 4 wheels, each labelled with ten digits i.e., from 0 to 9. The lock opens with a sequence of four digits with no repeats. What is the probability of a person getting the right sequence to open the suitcase?


A and B throw a pair of dice. If A throws 9, find B's chance of throwing a higher number.

 

Two unbiased dice are thrown. Find the probability that the sum of the numbers obtained on the two dice is neither a multiple of 2 nor a multiple of 3


A bag contains 8 red, 3 white and 9 blue balls. If three balls are drawn at random, determine the probability that all the three balls are blue balls 


A bag contains 8 red, 3 white and 9 blue balls. If three balls are drawn at random, determine the probability that  all the balls are of different colours.


A bag contains 5 red, 6 white and 7 black balls. Two balls are drawn at random. What is the probability that both balls are red or both are black?


Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(ii)
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]

In a single throw of three dice, find the probability of getting the same number on all the three dice.


A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability that all 10 are good


A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability thatat least one is defective


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is numbered 1, 2, 3, 4 or 5


Three squares of chessboard are selected at random. The probability of getting 2 squares of one colour and other of a different colour is ______.


If the letters of the word ALGORITHM are arranged at random in a row what is the probability the letters GOR must remain together as a unit?


Four candidates A, B, C, D have applied for the assignment to coach a school cricket team. If A is twice as likely to be selected as B, and B and C are given about the same chance of being selected, while C is twice as likely to be selected as D, what are the probabilities that A will not be selected?


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(A)


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(A ∪ B)


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(B ∩ C)


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all the three balls are red


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all one ball is red and two balls are white


If the letters of the word ASSASSINATION are arranged at random. Find the probability that no two A’s are coming together


In a non-leap year, the probability of having 53 tuesdays or 53 wednesdays is ______.


Without repetition of the numbers, four-digit numbers are formed with the numbers 0, 2, 3, 5. The probability of such a number divisible by 5 is ______.


A single letter is selected at random from the word ‘PROBABILITY’. The probability that it is a vowel is ______.


The probability that a person visiting a zoo will see the giraffee is 0.72, the probability that he will see the bears is 0.84 and the probability that he will see both is 0.52.


The probability that a student will pass his examination is 0.73, the probability of the student getting a compartment is 0.13, and the probability that the student will either pass or get compartment is 0.96.


The probabilities that a typist will make 0, 1, 2, 3, 4, 5 or more mistakes in typing a report are, respectively, 0.12, 0.25, 0.36, 0.14, 0.08, 0.11. 


The probability that the home team will win an upcoming football game is 0.77, the probability that it will tie the game is 0.08, and the probability that it will lose the game is ______.


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×